OpenStudy (anonymous):

prove the following identities: (a) tan(A-B)=(1-cot A tan B)/(cot A+tan B) (b) cot(A-B)=(cot A cat B+1)/(cot B-cot A)

OpenStudy (hba):

You know formula of tan(A-B)

OpenStudy (anonymous):

tan(A-B)=(tan A-tan B)/(1+tan A tan B)

OpenStudy (anonymous):

divide each term by tan A

OpenStudy (anonymous):

to prove the identity

OpenStudy (dumbcow):

just a matter of rearranging the fraction cotA = 1/tanA \[\large \frac{1-\frac{\tan B}{\tan A}}{\frac{1}{\tan A}+\tan B} = \frac{\frac{\tan A-\tan B}{\tan A}}{\frac{1+\tan A \tan B}{\tan A}} = \frac{\tan A -\tan B}{1+\tan A \tan B}=\tan(A-B)\]